OpenStudy (anonymous):

Linda owes $3,861 on a credit card with a 22.3% interest rate compounded monthly. What is the monthly payment she should make in order to pay off this debt in 2 years, assuming she does not charge any more purchases with the card? $160.88 $200.87 $361.93 $867.93

OpenStudy (amistre64):

arent you the one who said my Payment formula was "not textbook"?

OpenStudy (amistre64):

\[B_ok^{12*2}\frac{1-k}{1-k^{12*2}}=P\]

OpenStudy (amistre64):

such that k = 1+r/12

OpenStudy (anonymous):

No that wasn't me?

OpenStudy (amistre64):

lol :)